# Explanation User Interfaces: A Systematic Literature Review

### v1

@@ 8. The HERMES Platform to Guide Practitioners in Creating XUIs
This section presents a set of conclusions by relating the dimensions analyzed in our RQs. We examine how each dimension interacts across all RQs to identify recurring patterns. This approach allows us to distill insights that can support practitioners in navigating the broader landscape of the findings.
Visual explanations and text are the primary means of conveying explanations as shown in Figure 6, where the explanation modality is shown relative to its use within different domains, types of data, and types of users. Visual explanations are preponderant, being used more often than textual descriptions. This suggests that visual explanations may aid interpretability. This preference is particularly evident when working with tabular data and images, where graphical representations can make patterns and relationships more apparent. However, visual explanations are not limited to these data types; they are also applied to text and time series, while text-based explanations remain common for tabular data. It is also interesting to note how in the mobility domain (Schneider et al., 2021b), audio is present.
The choice of explanation technique often depends on the audience. SHAP (Malandri et al., 2023; Ma et al., 2022; Chromik, 2021; He et al., 2024; Jansen et al., 2024) explanations tend to be preferred when working with non-expert users due to their intuitive visualization of feature contributions. However, exemplars—i.e., concrete instances that illustrate a model’s behavior—and salient masks, which highlight the most influential input regions, are also commonly used to aid understanding (Huang et al., 2022; Wang and An, 2021; Yang et al., 2020; Larasati et al., 2023; Nakao et al., 2022).
When designing user studies for XUIs, the choice of method should align with the target user group. Interviews are most commonly used for domain experts, particularly in high-stakes fields like healthcare and finance, where qualitative insights are crucial for understanding decision-making needs. In contrast, surveys and usability studies are more prevalent among AI experts, emphasizing practicality and efficiency in technical contexts.
Regarding evaluation metrics, domain experts in high-stakes applications tend to prioritise helpfulness (Hwang et al., 2022b; Nourani et al., 2022b; Liu et al., 2022; Zytek et al., 2022) over transparency (Sevastjanova et al., 2021). Their primary concern is whether an explanation supports their decision-making process rather than simply revealing model internals. For XUIs designers, this means that explanations should be crafted to disclose system behaviour and integrate seamlessly into expert workflows, ensuring they provide meaningful, actionable insights.
8.1. Introducing the HERMES Platform
[…]
Figure 7. The HERMES Framework with the 6 dimensions observed in our analysis
[…]
@@ 8.1. Introducing the HERMES Platform
HERMES (a framework for Human-cEnteRed developMent of Explainable user interfaceS) is a web platform designed to support practitioners in building effective XUIs. Its name comes from Greek mythology, as Hermes was the messenger of the gods, and the god of language, representing the idea of delivering explanations efficiently and clearly. HERMES offers a structured collection of design guidelines derived from the SLR, presented as interactive cards that synthesize insights from the referenced papers. Each card includes a concise description of a guideline, relevant tags that reflect key dimensions from the associated paper, and a direct link to the original source. Users can explore the guidelines in two complementary ways: by browsing the full collection of cards, or by filtering recommendations through up to six entry points (fig. 7), which correspond to six of the thirteen dimensions of this survey (see table 2): (1) Application Domain, (2) User type, (3) AI model, (4) AI task, (5) XAI model, and (6) Explanation modality.
HERMES was developed using Jekyll [^6] https://jekyllrb.com/ and enables designers to either align their XUIs with an existing use context or to explore potential design directions under specific project constraints, fostering both conformity and creativity in the design process. Not all the input dimensions must be defined by the designer. Moreover, the output of the platform also includes data about the techniques to employ for the evaluation of the XUI (i.e., the Type of Study to conduct and the Metrics to measure). The platform is available at the following address: http://espositoandrea.github.io/hermes-mirror/. A screenshot of the web platform is reported in fig. 8.
@@ 8.2. Use case: a medical application
A medical technology company is developing a XUI for an existing AI-based system designed to assist clinicians in diagnosing cervical cancer using MRI scans. The existing software that currently helps clinicians is a neural network-based system that determines whether a patient’s blood test suggests that the clinical criteria for diabetes are met. However, the AI model only gives the clinician a binary outcome (yes/no), without producing an explanation. The development team is thus tasked to implement an XAI model and visualise the outputs for the clinicians, who need transparent and interpretable AI outputs to support their diagnostic decisions.
To guide the design of this XUI, the development team uses HERMES to determine the most suitable explanation and UI design strategies by setting four out of the six entry points as project constraints: Health as the application domain, Neural Network as the AI model, Image classification (cancerous vs. non-cancerous MRI scans) as the AI task, and Expert as the user type, as the interface is targeted to domain experts (i.e., clinicians). Since there are no constraints on the XAI technology to implement and how to show the results, the other two entry points (XAI model and Explanation modality) are left empty – thus becoming outputs of HERMES, which suggests what solutions to explore in this regard.
Based on the provided parameters, HERMES suggests explanation methods that have been effective in similar medical applications. The system highlights Counter-exemplars/factual, Decision Rules, Exemplars, Feature Importance, Neuron Activation, Salient Mask, Sensitivity Analysis, and Shapley Values as suitable XAI techniques. For explanation modalities, it recommends Natural Language, Text, and Visual explanations.
@@
One specific design guideline suggested by HERMES is Provide Contextual Information. This guideline states that placing explanations together with additional contextual information enhances their relevance and interpretability for domain-expert users, such as physicians. Moreover, the depth of information in explanations should be adaptable to the user’s needs, ensuring that both detailed and summary information is available without overwhelming them. To support this recommendation, HERMES provides references to relevant academic sources, including a specific paper by Cheng et al. (Cheng et al., 2022). In the paper, the developers can find more implementation details and access a visual example to guide the design of the new XUI. To ensure the XUI meets clinicians’ needs, HERMES suggests that user studies for designing and evaluating such interfaces should include interviews and user observation studies. This may be interpreted as a recommendation to conduct interviews with clinicians early in the design process, and then use observational studies to iteratively test and refine the XUI, following a human-centered design approach (ISO, 2019).
Using HERMES, the design team efficiently selects an adequate explanation approach tailored to expert clinicians in a medical domain. The platform informs the selection of XAI models and explanation modalities, and also suggests a human-centered approach to involve end-users both in the design and evaluation of the XUI. This structured methodology can indeed accelerate the design process, reducing trial-and-error efforts and ensuring that the final interface aligns with best practices found in the academic literature.
@@ 9. Future Challenges
Among the future challenges identified by this SLR, a key issue is the limited adoption of co-design practices and focus group methodologies in the development of HCXAI systems. These approaches are essential to strengthen a real co-creation dynamic between users and XAI systems, ensuring that user needs, mental models, and expectations are meaningfully integrated into the design process. Similarly, the collection of quantitative data and the execution of usability studies involving both AI and domain experts are critical for increasing the practical applicability and reliability of these tools.
Recent research acknowledges the growing role of Large Language Models in XAI. Lei et al. suggest that with the rising adoption of LLMs, the demand for XAI among non-expert users will increase. They propose a set of design principles for XUIs (Lei et al., 2024). LLMs are also beginning to be incorporated into Explainable Conversational Interfaces. For example, Joshi et al. conducted a Wizard of Oz user study comparing two versions of a vacation planning chatbot—one with low explainability and one with high explainability—involving 60 participants. Their study finds that providing explanations enhances trust and acceptance of the LLM-based system (Joshi et al., 2024), LLMs represent both a novel challenge and a promising opportunity for the development of XUIs. They are reshaping user-explanation interactions and advancing the potential for more conversational, human-like explanations, as envisioned by Miller (Miller, 2019). However, LLMs themselves remain opaque, prone to hallucinations, and can undermine user trust due to their lack of transparency and controllability. Addressing these matters is key for the future of effective and responsible XUIs design.
@@
10. Threats To Validity
In general, several threats to validity can undermine the results of a Systematic Literature Review. In the following section, we report the most common ones, detailing how we mitigated them.
  • Selection Bias. This occurs when the studies included in the review are not representative of the entire population of studies on the topic. This has been mitigated by: i) manually reviewing the publications to ensure their compliance with the SLR goal, and ii) performing two phases, i.e., search on digital libraries and snowballing. However, our search focused only on full publications; therefore, we acknowledge that this SLR does not cover demos, posters, and working papers presented at workshops.
  • Publication Bias. This occurs when studies that show statistically significant results are more likely to be published than studies that do not. This aspect has been mitigated by manually reading those publications that do not report any results but only a technical solution with preliminary results. Besides the generic inclusion criteria, their relevance for our SLR is considered, for example, the number of citations and the novelty of the solution.
  • Time Lag Bias. This occurs when the review does not include all relevant studies because they were published after the review was conducted. In this case, we can safely assume that this threat is not so evident in our study since the SLR was performed 5 months before its submission.
  • Publication Quality. This occurs when studies of poor quality are included in the review. To mitigate this aspect, we defined inclusion criteria on the quality of the venue of the publication, leaving to a manual evaluation of the authors of this SLR the inclusion of publications that appeared in venues of lower quality.
  • Grey Literature was not considered. This includes preprints and tools—such as commercial tools and platforms—that were not published in academic venues, despite their popularity or relevance to our RQs.
@@ 11. Conclusions
This study explores the design and evaluation of XUI. While XUIs are widely recognised as a fundamental component of the explanation process, their implementation still presents significant research challenges.
We approached this problem from a multidimensional perspective. On one hand, we examined HCI requirements, analyzing user types and evaluation metrics derived from user studies. On the other hand, we investigated algorithmic aspects, bridging these two dimensions through the interface layer.
Our analysis led to the development of guidelines for designers, combining both top-down and bottom-up approaches. The top-down perspective involved reviewing studies that propose directives and frameworks, while the bottom-up approach focused on extracting design principles and requirements from case studies. These guidelines aim to close the gap between the usability of explainable interfaces and approaches that focus solely on algorithmic transparency.
By providing a comprehensive view of the ecosystem in which XUIs are designed and implemented, this study helps identify key challenges in creating explainable processes that comply with emerging regulations. Ultimately, our findings contribute to making XAI not only a research concept but a practical and deployable solution across real-world applications. In fact, in this work, we also introduced HERMES, a platform for guiding practitioners in the design and evaluation of XUIs. Although user studies to evaluate and validate the platform are needed, with this platform, we attempt to bridge the gap that exists between results coming from the academic literature and their actual application by designers and developers.
@@
Conceptualization: EC; .
Acknowledgements.
This research is partially funded by (i) the Sponsor Italian Ministry of University and Research (MUR) and by the European Union - NextGenerationEU , Mission 4, Component 2, Investment 1.1, under grant PRIN 2022 “DevProDev: Profiling Software Developers for Developer-Centered Recommender Systems”—Grant #CUP: H53D23003620006; (ii) the Sponsor Italian Ministry of University and Research (MUR) and by the European Union - NextGenerationEU , Mission 4, Component 2, Investment 1.1, under grant PRIN 2022 PNRR “PROTECT: imPROving ciTizEn inClusiveness Through Conversational AI” Grant #(Grant P2022JJPBY)—CUP: H53D23008150001. (iii) the co-funding of the Sponsor European Union—Next Generation EU : NRRP Initiative, Mission 4, Component 2, Investment 1.3 – Partnerships extended to universities, research centres, companies, and research D.D. MUR n. 341 del 15.03.2022 – Next Generation EU (PE0000013 – “Future Artificial Intelligence Research - FAIR” - Spoke 6 “Symbiotic AI”—Grant #CUP: H97G22000210007). (iv) the Partnership Extended Grant #PE00000013 - Sponsor “FAIR - Future Artificial Intelligence Research” - Spoke 1 “Human-centered AI” (vi) Sponsor TANGO project , grant agreement no. Grant #101120763, (vii) The Sponsor European Community Horizon 2020 programme under the funding scheme Grant #ERC-2018-ADG G.A. 834756 XAI: Science and technology for the eXplanation of AI decision making The research of Andrea Esposito is funded by a Ph.D. fellowship within the framework of the Sponsor Italian “D.M. n. 352, April 9, 2022” - under the National Recovery and Resilience Plan, Mission 4, Component 2, Investment 3.3 - Ph.D. Project “Human-Centered Artificial Intelligence (HCAI) techniques for supporting end users interacting with AI systems,” co-supported by “Eusoft S.r.l.” (Grant #CUP H91I22000410007). The research of Francesco Greco is funded by a Ph.D. fellowship within the framework of the Sponsor Italian “D.M. n. 352, April 9, 2022” - under the National Recovery and Resilience Plan, Mission 4, Component 2, Investment 3.3 - Ph.D. Project “Investigating XAI techniques to help users defend from phishing attacks”, co-supported by “Auriga S.p.A.” (Grant #CUP H91I22000410007).

### v2

@@ 5.1. Navigating the Results: Introducing the HERMES Platform
Figure 8. The HERMES Framework with the 6 dimensions observed in our analysis.
@@
HERMES (a framework for Human-cEnteRed developMent of Explainable user interfaceS) is a web platform designed to support practitioners in building effective XUIs. It offers a structured collection of design guidelines derived from the results of the SLR, presented as interactive cards that synthesize insights from the referenced papers. Each card includes a concise description of a guideline, relevant tags that reflect key dimensions from the associated paper, and a direct link to the original source. Users can explore the guidelines in two complementary ways: by browsing the full collection of cards, or by filtering them through up to six entry points (Fig. 8), which correspond to six of the thirteen dimensions of this survey (see Table 2): (1) Application Domain, (2) User type, (3) AI model, (4) AI task, (5) XAI model, and (6) Explanation modality. HERMES was developed using Jekyll [^6] https://jekyllrb.com/ and enables designers to either align their XUIs with an existing use context (by analyzing the above dimensions) or to explore potential design directions under specific project constraints, fostering both conformity and creativity in the design process. Not all the input dimensions must be defined by the designer. Moreover, the output of the platform also includes data about the techniques to employ for the evaluation of the XUI (i.e., the Type of Study to conduct and the evaluation constructs to assess). The platform is available at the following address: http://espositoandrea.github.io/hermes-mirror/. A screenshot of the web platform is reported in Fig. 9.
@@ 5.2. Use case: a medical application
Let us suppose that a medical technology company is developing a XUI for an existing AI-based system designed to assist clinicians in diagnosing cervical cancer using MRI scans. The existing software that currently helps clinicians is a neural network-based system that determines whether a patient’s blood test suggests that the clinical criteria for diabetes are met. However, the AI model only gives the clinician a binary outcome (yes/no), without providing an explanation. The development team is thus tasked to implement an XAI model and visualise the outputs for clinicians, who need transparent and interpretable AI outputs to support their diagnostic decisions.
@@
To guide the design of this XUI, the development team uses HERMES to determine the most suitable explanation and UI design strategies by setting four out of the six entry points as project constraints: Health as the application domain, Neural Network as the AI model, Image classification (cancerous vs. non-cancerous MRI scans) as the AI task, and Domain Expert as the user type, as the interface is targeted to domain experts (i.e., clinicians). Since there are no other design constraints, the other two entry points (XAI model and Explanation modality) are left empty – thus becoming suggestions from HERMES. Based on the provided parameters, HERMES suggests explanation methods that have been effective in similar medical applications. The system highlights several XAI techniques as suitable, including Counter-exemplars/factual and Decision Rules, and Natural Language, Text, and Visual explanations as explanation modalities. HERMES also identifies the design guideline Provide Contextual Information, which states that placing explanations alongside additional contextual information enhances their relevance and interpretability for domain-expert users, such as physicians, while tailoring the depth of explanations to the user’s needs. To support this recommendation, HERMES provides references to relevant academic sources, including a specific paper by Cheng et al. (2022). In the paper, the developers can find more implementation details and access a visual example to guide the design of the new XUI. To ensure the XUI meets clinicians’ needs, HERMES suggests that user studies for designing and evaluating such interfaces should include interviews and user observation studies. This may be interpreted as a recommendation to conduct interviews with clinicians early in the design process, and then use observational studies to iteratively test and refine the XUI, following a human-centered design approach (ISO, 2019b). Using HERMES, the design team efficiently selects an adequate explanation approach tailored to expert clinicians in a medical domain. The platform informs the selection of XAI models and explanation modalities, and also suggests a human-centered approach to involve end-users both in the design and evaluation of the XUI. This structured methodology can indeed accelerate the design process, reducing trial-and-error efforts and ensuring that the final interface aligns with best practices found in the academic literature.
@@ 6. Future Challenges
Among the future challenges identified by this SLR, a key issue is the limited adoption of co-design practices and focus group methodologies in the development of HCXAI systems. These approaches are essential to strengthen a real co-creation dynamic between users and XAI systems, ensuring that user needs, mental models, and expectations are meaningfully integrated into the design process. Similarly, the collection of quantitative data and the execution of usability studies involving both AI and domain experts are critical for increasing the practical applicability and reliability of these tools. Recent research acknowledges the growing role of Large Language Models in XAI. Lei et al. suggest that with the rising adoption of LLMs, the demand for XAI among non-expert users will increase. They propose a set of design principles for XUIs (Lei et al., 2024). LLMs are also beginning to be incorporated into Explainable Conversational Interfaces. For example, Joshi et al. conducted a Wizard of Oz user study comparing two versions of a vacation planning chatbot—one with low explainability and one with high explainability—involving 60 participants. Their study finds that providing explanations enhances trust and acceptance of the LLM-based system (Joshi et al., 2024), LLMs represent both a novel challenge and a promising opportunity for the development of XUIs. They are reshaping user-explanation interactions and advancing the potential for more conversational, human-like explanations, as envisioned by Miller (Miller, 2019). However, LLMs themselves remain opaque, prone to hallucinations, and can undermine user trust due to their lack of transparency and controllability. Addressing these matters is key for the future of effective and responsible XUIs design.
@@
7. Limitations and Threats To Validity
In general, several threats to validity can undermine the results of a Systematic Literature Review. In the following section, we report the most common ones, detailing how we mitigated them. Selection Bias: This occurs when the studies included in the review are not representative of the entire population of studies on the topic. This has been mitigated by: i) manually reviewing the publications to ensure their compliance with the SLR goal, and ii) performing two phases, i.e., search on digital libraries and snowballing. However, our search focused only on full publications; therefore, we acknowledge that this SLR does not cover demos, posters, and working papers presented at workshops. Publication Bias: This occurs when studies that show statistically significant results are more likely to be published than studies that do not. This aspect has been mitigated by manually reading those publications that do not report any results but only a technical solution with preliminary results. Besides the generic inclusion criteria, their relevance for our SLR is considered, for example, the number of citations and the novelty of the solution. Time Lag Bias: This occurs when the review does not include all relevant studies because they were published after the review was conducted. In this case, we can safely assume that this threat is not so evident in our study since the SLR was performed 5 months before its submission. Publication Quality: This occurs when studies of poor quality are included in the review. To mitigate this aspect, we defined inclusion criteria on the quality of the venue of the publication, leaving to a manual evaluation of the authors of this SLR the inclusion of publications that appeared in venues of lower quality. Grey Literature was not Considered: This includes preprints and tools, such as commercial tools and platforms, that were not published in academic venues, despite their popularity or relevance to our RQs.
The organization and navigation criteria adopted in this SLR should not be interpreted as exhaustive or definitive since they come from the authors’ design choices. Alternative classification schemes or additional dimensions may be equally valid and could offer complementary insights, particularly for domain-specific or practice-oriented use cases. Moreover, the effectiveness and appropriateness of these navigation criteria have not yet been empirically evaluated with end users. To mitigate this limitation and enable stakeholders to customize their navigation and obtain insights into our results, we provide access to HERMES. Additionally, while we acknowledge the importance of user-based validation of HERMES, we consider it as a supporting and exploratory artifact within the broader scope of this work—which synthesizes and structures existing knowledge on XUIs—rather than as a standalone, fully validated system. HERMES was in fact designed to complement this contribution by offering an interactive and actionable way to navigate the results of the SLR, facilitating exploration by researchers, designers, and practitioners. A comprehensive user-centered evaluation of HERMES (in particular assessing usability, usefulness, and impact on research and design practices) is therefore left as an important direction for future work. Finally, findings were reported by clusters to maintain contextual specificity and avoid overgeneralization across multiple domain contexts. This approach indeed limits the interpretational boundaries of the results since cross-domain transferability would require dedicated user studies, thus strengthening the methodological rigor of the SLR and preventing misleading conclusions.
@@ 8. Conclusions
This study explores the design and evaluation of XUI. While XUIs are widely recognised as a fundamental component of the explanation process, their implementation still presents significant research challenges. We approached this problem from a multidimensional perspective. On one hand, we examined HCI requirements, analyzing user types and evaluation constructs derived from user studies. On the other hand, we investigated algorithmic aspects, bridging these two dimensions through the interface layer. Our analysis led to the development of guidelines for designers, combining both top-down and bottom-up approaches. The top-down perspective involved reviewing studies that propose directives and frameworks, while the bottom-up approach focused on extracting design principles and requirements from case studies. These guidelines aim to close the gap between the usability of explainable interfaces and approaches that focus solely on algorithmic transparency. By providing a comprehensive view of the ecosystem in which XUIs are designed and implemented, this study helps identify key challenges in creating explainable processes that comply with emerging regulations. Ultimately, our findings contribute to making XAI not only a research concept but a practical and deployable solution across real-world applications. In fact, in this work, we also introduced HERMES, a platform for guiding practitioners in the design and evaluation of XUIs. Although user studies to evaluate and validate the platform are needed, with this platform, we attempt to bridge the gap that exists between results coming from the academic literature and their actual application by designers and developers.